OpenStudy (anonymous):

Find the equation of the circle with center at (3, 2) and through the point (5, 4)

OpenStudy (anonymous):

Can you work out the radius of the circle from what you're given?

OpenStudy (anonymous):

I don't know how :/

OpenStudy (anonymous):

Ok so you want the distance between two points, it's a formula \[distance = \sqrt{ (x_2 - x_1)^2 + (y_2 - y_1)^2} \] You're x2 and y2 are the coordinates of (5,4) and x1 and y1 are the coordinates of (3,2) :) Now can you work out the distance?

OpenStudy (anonymous):

Ok, I tried. I got 1.41 but I'm not sure if I did it correctly

OpenStudy (anonymous):

Ok so I did \[\sqrt{(5-3)^2 + (4-2)^2} = 2\sqrt{2} = 2.824\]

OpenStudy (anonymous):

Ok, what next?

OpenStudy (anonymous):

I worked it out and got 8 as an answer :) Thanks for the help!